OpenStudy (anonymous):

Guestion: x − 4y + 2z − 3x − 3y + 2z = ? −2x + 7y + 4z −2x −7y + 4z −3x + 7y + 4z −3y − 7z + 4z

OpenStudy (anonymous):

\[ \left( 1 -3 \right) x + \left( -4 -3 \right) y + \left( 2 + 2 \right) z = -2x -7y +4z \]